### v1

# Knapsack Optimization-based Schema Linking for LLM-based Text-to-SQL Generation

Abstract
Generating SQLs from user queries is a long-standing challenge, where the accuracy of initial schema linking significantly impacts subsequent SQL generation performance. However, current schema linking models still struggle with missing relevant schema elements or an excess of redundant ones. A crucial reason for this is that commonly used metrics, recall and precision, fail to capture relevant element missing and thus cannot reflect actual schema linking performance. Motivated by this, we propose an enhanced schema linking metric by introducing a restricted missing indicator. Accordingly, we introduce Knapsack optimization-based Schema Linking Agent (KaSLA), a plug-in schema linking agent designed to prevent the missing of relevant schema elements while minimizing the inclusion of redundant ones. KaSLA employs a hierarchical linking strategy that first identifies the optimal table linking and subsequently links columns within the selected table to reduce linking candidate space. In each linking process, it utilize a knapsack optimization approach to link potentially relevant elements while accounting for a limited tolerance of potential redundant ones. With this optimization, KaSLA-1.6B achieves superior schema linking results compared to large-scale LLMs, including deepseek-v3 with state-of-the-art (SOTA) schema linking method. Extensive experiments on Spider and BIRD benchmarks verify that KaSLA can significantly improve the SQL generation performance of SOTA text-to-SQL models by substituting their schema linking processes.
Knapsack Optimization-based Schema Linking for
LLM-based Text-to-SQL Generation
Zheng Yuan1, Hao Chen2, Zijin Hong1, Qinggang Zhang1 Feiran Huang3, Xiao Huang1 1The Hong Kong Polytechnic University 2City University of Macau, 3Jinan University {yzheng.yuan, zijin.hong, qinggangg.zhang}@connect.polyu.hk sundaychenhao@gmail.com; huangfr@jnu.edu.cn; xiao.huang@polyu.edu.hk
1 Introduction
Figure 1: In text-to-SQL tasks, commonly used schema linking metrics such as Recall and Precision fail to accurately capture the true performance of schema linking when relevant schema elements are missing.
Figure 2: Overall framework of proposed knapsack optimization-based schema linking agent (KaSLA) with column linking as an example. Traditional schema linking models often struggle with both missing relevant elements and the inclusion of redundant ones. KaSLA addresses this challenge by maximizing the total estimated relevance while maintaining total redundancy within a specified tolerance using knapsack optimization. Serving as a general model, KaSLA enhances the SQL generation of text-to-SQL models by replacing their existing schema linking processes.
With the advent of large language models (LLMs) (Achiam et al., 2023; Dubey et al., 2024), LLM-based text-to-SQL is emerging as the next-generation interface for database users (Hong et al., 2024b). Typically, text-to-SQL frameworks employ a two-step process: first linking database schema elements (tables and columns) to user queries, then generating the corresponding SQL statements. However, the accurate SQL generation requires the linking to the correct schema elements, meaning that errors in early stages are inevitably propagated to later SQL generation. For instance, inaccurate schema linking—such as overlooking relevant schema elements or linking to irrelevant ones—will lead to incorrect SQL statements.
Recent state-of-the-art text-to-SQL models tend to focus primarily on final SQL generation while employing relatively simplistic schema linking strategies or neglecting it. DTS-SQL (Pourreza and Rafiei, 2024) finetunes a local LLM to select relevant tables based on user queries. Solid-SQL (Liu et al., 2025) present a data augmentation method for the finetuning of Schema linking method. Chess (Talaei et al., 2024) and CodeS (Li et al., 2024b) use semantic-based retrieval models to enhance the schema linking process by retrieving relevant information from databases. RSL-SQL (Cao et al., 2024) introduces a bidirectional approach to simplify the full schema. E-SQL (Qu et al., 2024) and (Maamari et al., 2024) assert that schema linking remains challenging for current LLMs and thus directly use the full schema as input for SQL generation.
Even though schema linking is crucial, current widely used metrics, such as Precision and Recall (Li et al., 2023a, 2024b; Qu et al., 2024), cannot adequately evaluate instances when relevant elements are missing. As demonstrated in Figure 1, missing a single relevant table or column can significantly decrease the accuracy of SQL generation, yet still result in high Recall and Precision scores. Because Recall does not significantly penalize the model as long as the majority of the relevant elements are identified, allowing it to achieve an artificially high score, even when crucial elements are missing. Similarly, regardless of the severity of the missing, as long as the number of correctly linked elements is proportionally high compared to the total predicted relevant elements, the model can achieve an unreasonably high Precision score. This discrepancy distorts the assessment of a schema linking model’s effectiveness, potentially hindering progress in text-to-SQL development.
To address this, we firstly design a tailored schema linking metrics that accurately assess linking performance with a restricted element missing indicator. Additionally, we propose the Knapsack optimization-based Schema Linking Agent (KaSLA), designed to link the most relevant and minimally redundant schema elements according to the given queries. KaSLA introduces a binary-probabilistic score function to predict a robust relevance score of each element. Additionally, KaSLA identifies redundancy of each element and the upper redundancy tolerance for each query based on its most similar queries in the training dataset, effectively minimizing the inclusion of redundant elements. Finally, KaSLA links tables and columns with an hierarchical strategy to reduce the column linking candidate space for efficiency and employs a 0-1 knapsack optimization approach to select the most relevant element set under the upper redundancy tolerance to reduce the redundant linking. Extensive experiments demonstrate that KaSLA can improve SQL generation accuracy by reducing missing and redundant information in the schema linking.
In summary, our contributions are as follows:
  • We introduce an enhanced schema linking metric, which can accurately capture the missing of relevant elements and reflect the actual linking performance aligning with the SQL generation performance.
  • We propose the Knapsack optimization-based Schema Linking Agent (KaSLA) to ensure the linking of relevant schema elements and minimizing the inclusion of redundant ones.
  • KaSLA can serve as a general agent to enhance LLM-based text-to-SQL framework by provide optimal linking results. Extensive experiments demonstrate improved performance in the SOTA text-to-SQL models when using KaSLA for schema linking.
2 Preliminaries
In this section, we outline the definitions for the schema linking and 0-1 Knapsack Optimization.
Schema Linking Task.
Given a natural language query $q$ and a corresponding database schema ${\mathcal{S}}_{q}$ containing schema elements, tables ${\mathcal{T}}=\{t_{1},\cdots,t_{|{\mathcal{T}}|}\}$ and columns ${\mathcal{C}}=\{c_{1},\cdots,c_{|{\mathcal{C}}|}\}$ in each table. Schema linking is a common process used in text-to-SQL to predict the relevant schema entities with $q$ from ${\mathcal{S}}_{q}$:
$f_{\text{linking}}:(q,{\mathcal{S}}_{q})\mapsto\widehat{{\bm{S}}}_{q},$ (1)
where $\widehat{{\bm{S}}}_{q}$ is the subset of ${\mathcal{S}}_{q}$ that includes schema element $s\in{\mathcal{S}}_{q}$ that relevant to $q$. Each $s$ could represent a specific table or column. Additionally, we denote ${\bm{S}}_{q}^{*}$ as the ground truth linking results.
0-1 Knapsack Optimization.
The 0-1 knapsack problem Fréville (2004) is a classic combinatorial optimization challenge. It involves selecting items from a finite item set, each with a value ${\mathcal{V}}$ and weight ${\mathcal{W}}$, to place into a knapsack with a maximum weight capacity ${\mathcal{C}}$. Each item can be selected only once. Formally, the objective of 0-1 knapsack optimization is to select items such that their total weight remains below ${\mathcal{C}}$ while maximizing total value of these selected items:
$\widehat{N}=\arg\max\sum_{i\in\widehat{N}}{\mathcal{V}}_{i},\ \text{s.t.}\sum_{i\in\widehat{N}}{\mathcal{W}}_{i}\leq{\mathcal{C}},$ (2)
where $\widehat{N}$ is the selected item set for the knapsack.
3 Enhanced Schema Linking Metrics
In this section, we introduce enhanced schema linking metrics considering the missing of relevant elements during evaluation, thereby accurately assess the true performance of schema linking models.
For a query $q$ with predicted schema linking result $\widehat{{\bm{S}}}_{q}$ and ground truth linking result ${\bm{S}}^{*}_{q}$, Recall and Precision are two commonly used metreics (Li et al., 2023a, 2024b; Qu et al., 2024). Recall, ${\bm{R}}(q)=\frac{|\widehat{{\bm{S}}}_{q}\cap{\bm{S}}^{*}_{q}|}{|{\bm{S}}^{*}_{q}|}$, measures the proportion of relevant elements correctly identified, specifically how many of the actual relevant elements in ${\bm{S}}^{*}_{q}$ are found in $\widehat{{\bm{S}}}_{q}$. Conversely, Precision, ${\bm{P}}(q)=\frac{|\widehat{{\bm{S}}}_{q}\cap{\bm{S}}^{*}_{q}|}{|\widehat{{\bm{S}}}_{q}|}$, assesses the accuracy of predictions by determining the proportion of elements predicted as relevant in $\widehat{{\bm{S}}}_{q}$ that are indeed correct.
As illustrated in Figure 1, these metrics can yield high scores even in the missing of relevant elements. This underscores that they do not effectively address the issue of missing relevant elements, which can undermine the true accuracy of schema linking models.
To address the limitations of Recall and Precision, we firstly introduce a restricted element missing indicator to detect whether relevant elements are missing in the predicted schema linking result $\widehat{{\bm{S}}}_{q}$, The formal definition is as follows:
$\bm{\mathbbm{1}}(q)=\begin{cases}1,&\text{if }\widehat{{\bm{S}}}_{q}\cap{\bm{S}}^{*}_{q}={\bm{S}}^{*}_{q},\\ 0,&\text{otherwise}.\end{cases}$ (3)
It takes a value of 1 when the predicted result $\widehat{{\bm{S}}}_{q}$ contain all the elements in the ground truth ${\bm{S}}^{*}_{q}$, meaning no relevant elements are missing. Conversely, it is set to 0 when there are any missing. Building on this missing indicator, we define enhanced Recall and Precision as follows:
${\bm{R}}^{\text{{+}}}(q)=\bm{\mathbbm{1}}(q){\bm{R}}(q),{\bm{P}}^{\text{{+}}}(q)=\bm{\mathbbm{1}}(q){\bm{P}}(q).$ (4)
Such metrics strengthen the penalty for missing relevant elements, as schema linking results with missing are inherently incorrect for SQL generation. For simplicity, we use the harmonic mean of enhanced Recall and Precision, referred to as the enhanced F1 score, as the overall measure for evaluating schema linking results:
${{\bm{F}}1}^{\text{{+}}}(q)=\frac{2{\bm{R}}^{\text{{+}}}(q){\bm{P}}^{\text{{+}}}(q)}{{\bm{R}}^{\text{{+}}}(q)+{\bm{P}}^{\text{{+}}}(q)}.$ (5)
This approach provides a comprehensive evaluation of both recall and precision while incorporating the critical aspect of missing elements.
4 Knapsack Optimization-based Schema Linking Agent (KaSLA)
In this section, we introduce the proposed KaSLA framework, which focuses on ensuring the linking of relevant schema elements to avoid missing and minimizing the inclusion of redundant ones. As illustrated in Figure 2, We begin by detailing a binary-probabilistic score function used for assessing the relevance of schema elements and an estimation function for the redundancy tolerance of each schema linking process. We then outline the knapsack optimization approach to link potential relevant element while reducing the potential redundant ones under the tolerance constraint. Finally, we describe the hierarchical schema linking process, demonstrating its efficiency for reducing the column linking candidate.
4.1 Relevance Estimation
To precisely assess the relevance of each schema element $s$ in relation to the given query $q$, we employ a hybrid binary-probabilistic score function. This function integrates a generative model for binary scoring alongside an encoding model for probabilistic scoring, allowing for a comprehensive estimation of relevance.
Binary scoring function.
This component is tasked with providing binary relevance assessments, determining whether a schema element is relevant or not in a straightforward manner:
$f_{\text{binary}}:(q,{\mathcal{S}}_{q})\mapsto\{r^{b}_{1},\cdots,r^{b}_{|{\mathcal{S}}_{q}|}\},$ (6)
where $r^{b}_{i}\in\{0,1\}$ represents the binary score assigned to the $i$-th element $s_{i}\in{\mathcal{S}}_{q}$. A score of 1 indicates relevant while 0 signifies redundant.
Considering efficiency, we finetune a lightweight LLM, DeepSeek-Coder-1.3B (Guo et al., 2024), using LoRA (Hu et al., 2021), a well-established parameter-efficient technique. We employ a strategy whereby the initial generation of a simulated SQL provides a contextual foundation for the subsequent generation of schema linking. The loss function during fine-tuning is defined as follows:
$\mathcal{L}_{\text{binary}}=-\sum_{q\in Q}\log P({\text{SQL}}^{*}_{q},{\bm{S}}_{q}^{*}\mid q,{\mathcal{S}}_{q}),$ (7)
where $Q$ is the set of queries in the training dataset. ${\text{SQL}}^{*}_{q}$ and ${\bm{S}}^{*}$ are the ground truth SQL and linking result, respectively. This formulation utilizes a joint generation of a simulated SQL and the schema linking result, thereby ensuring that the linking generation benefits from the contextual alignment provided by the simulated SQL during both fine-tuning and inference. Noted that the simulated SQL generated during this process will not be utilized in subsequent steps. The input and output examples are provided in Table 6 in Appendix B.
Probabilistic scoring function.
In conjunction with the binary scoring model, we introduce a probabilistic scoring model to enhance prediction robustness. The primary motivation is to mitigate the risk of the binary model misclassifing relevant elements. The probabilistic scoring model assigns a soft score to each element as the linking probability to reduce the likelihood of relevant elements missing , thereby supporting the binary model:
$f_{\text{prob.}}:(q,{\mathcal{S}}_{q})\mapsto\{r^{p}_{1},\cdots,r^{p}_{|{\mathcal{S}}_{q}|}\},$ (8)
where $r^{p}_{i}\in[0,1]$ represents the probabilistic score assigned to the $i$-th element $s_{i}\in{\mathcal{S}}_{q}$.
For the probabilistic scoring model, we utilize an encoding model, RoBERTa-Large (Liu et al., 2019), to derive the semantic embeddings for the query and all schema elements. Following the methodology of (Li et al., 2023a, 2024b), we employ a cross-attention network to jointly embed the semantic representations of tables and their columns. The dot product of the query embedding and the element embedding yields the final probabilistic score. This probabilistic model is trained using the presence of $s$ in ${\bm{S}}^{*}$ as the learning objective. The loss function is defined as follows:
$\mathcal{L}_{\text{prob.}}=\sum_{q\in Q}\text{FL}(\mathbbm{1}(s\in{\bm{S}}^{*}_{q}),f_{\text{prob.}}(q,{\mathcal{S}}_{q})),$ (9)
where $\text{FL}(\cdot)$ denotes the focal loss function (Ross and Dollár, 2017), designed to emphasize learning from hard negative samples.
Relevance scoring estimation.
The final function for estimating the relevance of each schema element integrates the scores derived from both the binary and probabilistic functions:
$\begin{split}f_{\text{relevance}}:&(q,{\mathcal{S}}_{q})\mapsto\{r_{1},\ldots,r_{|{\mathcal{S}}_{q}|}\}\\ &=\left\{\min(1,r^{b}_{i}+r^{p}_{i})\mid s_{i}\in{\mathcal{S}}_{q}\right\},\end{split}$ (10)
where $r_{i}$ denotes the relevance score assigned to the $i$-th schema element $s_{i}\in{\mathcal{S}}_{q}$ given the query $q$. An upper limit of 1 is imposed on $r_{i}$ to prevent excessively strong contrasts between potentially relevant elements, as each relevant schema element is equally important for accurate SQL generation.
4.2 Redundancy Estimation
To minimize the inclusion of redundant elements during schema linking, we define both a redundancy score for each schema element $s\in{\mathcal{S}}_{q}$ and the upper redundancy tolerance for query $q$.
Redundancy scoring estimation.
Recognizing that elements with lower relevance scores are more likely to be redundant, whereas those with higher scores are more essential, we assign a redundancy score $w_{i}$ to each schema element $s_{i}\in{\mathcal{S}}_{q}$ in relation to query $q$. This score is calculated as the inverse of its relevance score $r_{i}$, formulated as follows:
$w_{i}={r_{i}}^{-1}.$ (11)
It ensures elements with lower relevance scores contribute more to redundancy, thereby effectively prioritizing more relevant schema elements.
Upper Redundancy Tolerance.
Redundancy tolerance defines the limit on the total redundancy score of elements in the schema linking results, as an excess of redundant elements can significantly disrupt subsequent SQL generation. Estimating it during inference is challenging. For the SQL generation of some queries, a lenient tolerance may be beneficial since it allowing the inclusion of more potentially relevant elements in the linking results to prevent missing. However, for other queries, a lenient tolerance could introduce to many redundant elements, which may significantly disrupt the final SQL generation process.
To address this challenge, we hypothesize that similar queries exhibit comparable tolerance levels for redundancy and leverage queries from the training dataset, where ground truth linking results are available, to aid in predicting redundancy tolerance for given query $q$ during inference. We first retrieve the top-$K$ queries in the training dataset that closely resemble $q$ using sentence similarity measures (Gao et al., 2021). For a query $q_{k}$ in the retrieved top-$K$ queries, we calculate the total redundancy score for its ground truth linking results ${\mathcal{S}}^{*}_{q_{k}}$ and use it as its upper redundancy tolerance. The maximum of these upper redundancy tolerances from the $K$ most similar queries in the training dataset is then used as the redundancy tolerance estimate for query $q$:
$u_{q}=\max_{q_{k}\in\text{TopK}(q)}\bigg{(}\sum_{s_{j}\in{\mathcal{S}}^{*}_{q_{k}}}w_{j}\bigg{)}$ (12)
By capturing the maximum sum from the most similar queries, this approach effectively defines an upper bound on redundancy tolerance that is both efficient and adaptable, ensuring that while redundant elements are excluded, the risk of missing relevant elements is minimized.
4.3 Schema Linking with Knapsack Optimization
Given the relevance score $r_{i}$, the redundancy score $w_{i}$ for each $s_{i}\in{\mathcal{S}}_{q}$, and the upper redundancy tolerance $u_{q}$ for query $q$, we introduce a hierarchical schema linking strategy with knapsack optimization. This approach is designed to ensure the linking of relevant schema elements to avoid missing and minimize the inclusion of redundant ones for a precise schema linking.
Knapsack Optimization
The 0-1 knapsack problem (Fréville, 2004) is a classic combinatorial optimization challenge. It involves selecting items from a finite set, each item characterized by a specific value and weight, to place into a knapsack with a limited weight capacity. The objective of 0-1 knapsack optimization is to choose items that maximize the total value while ensuring the total weight remains below the knapsack’s capacity.
Inspired by this, we develop a knapsack optimization-based schema linking strategy to ensure the inclusion of relevant schema elements while minimizing redundant ones, thus ach ieving optimal schema linking. Formally, the optimization objective are defined as follows:
$\begin{split}f_{\text{Knap}}:&(q,{\mathcal{S}}_{q})\mapsto\widehat{{\bm{S}}}_{q}=\arg\max\sum_{s_{i}\in\widehat{{\bm{S}}}_{q}}r_{i}\\ &\text{s.t.}\quad\sum_{s_{i}\in\widehat{{\bm{S}}}_{q}}w_{i}\leq u_{q}.\end{split}$ (13)
This formulation aims to maximize the total relevance scores of the linked schema elements. The constraint ensures that the cumulative redundancy score remains within the upper redundancy tolerance of query $q$. To tackle this optimization problem and achieve the optimal linking result, We employ an efficient dynamic programming approach. Detailed pseudocode of this dynamic programming algorithm is provided in Algorithm A in Appendix.
Hierarchical Linking Strategy
For efficiency, our KaSLA employs a hierarchical process in schema linking: initially linking tables and subsequently linking columns within each selected table. This strategy effectively reduces the dimensionality of candidate columns, enhancing KaSLA’s capacity to handle large-scale schema linking and text-to-SQL generation tasks.
In the initial phase for a given query $q$, KaSLA aims to identify the relevant tables from the set of all available tables in schema ${\mathcal{S}}_{q}$, denoted as ${\mathcal{T}}_{q}$:
$\widehat{{\bm{T}}}_{q}=f_{\text{Knap}}(q,{\mathcal{T}}_{q}).$ (14)
After determining tables, for each $t\in\widehat{{\bm{T}}}_{q}$, let ${\mathcal{C}}^{t}_{q}$ represent the columns in table $t$. KaSLA then optimizes column linking within each selected table:
$\widehat{{\bm{C}}}^{t}_{q}=f_{\text{Knap}}(q,{\mathcal{C}}^{t}_{q}).$ (15)
The final schema linking result $\widehat{{\bm{S}}}_{q}$ is the union of all selected tables and their corresponding selected columns, represented as:
$\widehat{{\bm{S}}}_{q}=\widehat{{\bm{T}}}_{q}\cup\bigcup_{t\in\widehat{{\bm{T}}}_{q}}\widehat{{\bm{C}}}^{t}_{q}.$ (16)
We provide a complexity analysis is detailed in Appendix D, offering insights into KaSLA’s implementation and computational efficiency.
5 Experiments
We conducted comprehensive experiments to evaluate KaSLA and address the following research questions: RQ1: Does KaSLA enhance existing text-to-SQL baselines by improving their schema linking capabilities? RQ2: Does KaSLA demonstrate superior schema linking performance by handling element missing and redundancy? RQ3: Does each component in KaSLA contribute to the overall performance, and is KaSLA an efficient model for real-world text-to-SQL applications?
5.1 Experiment Setup
We conducted experiments on two well-know large-scale text-to-SQL datasets, BIRD (Li et al., 2023b) and Spider (Yu et al., 2018), incorporating multiple robust baselines based on LLMs with in-context learning and fine-tuning. For evaluating text-to-SQL, we utilized Execution Accuracy (EX) (Yu et al., 2018; Li et al., 2023b) and Valid Efficiency Score (VES) (Li et al., 2023b). We included multiple robust baselines from two categories:(i) LLMs + in-context learning baselines: C3-SQL (Dong et al., 2023), DIN-SQL (Pourreza and Rafiei, 2023), MAC-SQL (Wang et al., 2024), DAIL-SQL (Gao et al., 2024), SuperSQL (Li et al., 2024a), Dubo-SQL (Thorpe et al., 2024), TA-SQL (Qu et al., 2024), E-SQL (Caferoğlu and Ulusoy, 2024), CHESS (Talaei et al., 2024); and (ii) LLMs + fine-tuning baselines: DTS-SQL (Pourreza and Rafiei, 2024), CodeS (Li et al., 2024b).
5.2 Main Results of SQL Generation (RQ1)
To address RQ1, we conducted experiments to evaluate the SQL generation capabilities of state-of-the-art (SOTA) baselines enhanced by our KaSLA, with results presented in Table 1. KaSLA replaces the original schema linking component in models that already have schema linking capabilities or implements a full schema prompt substitution for those that do not.
The results demonstrate that incorporating KaSLA consistently improves both Execution Accuracy (EX) and Valid Efficiency Score (VES) across baselines with diverse base models, including GPT-4, CodeS-15B, StarCoder2-15B, and Deepseek-v3. Notably, in the BIRD-dev dataset, KaSLA enhances the EX of CHESS with Deepseek-v3 from 61.34% to 63.30%. For local LLMs, both "CodeS + KaSLA" and "TA-SL + KaSLA" with CodeS-15B and StarCoder-15B achieve impressive increases in Execution Accuracy compared to their original versions. These results underscore KaSLA’s effectiveness and flexibility as a plug-in, showing that KaSLA plays an important role in improving complex text-to-SQL tasks. This adaptability not only confirms its robustness but also broadens its applicability across different models, achieving superior performance outcomes and making it a versatile addition to existing systems.
Base Model Method BIRD-dev Spider-dev
EX VES EX VES
Easy Medium Hard Total Total Easy Medium Hard Extra Total Total
GPT-4 MAC-SQL - - - 57.56 58.76 - - - - 86.75 -
DIN-SQL - - - 50.72 58.79 92.34 87.44 76.44 62.65 82.79 81.70
DAIL-SQL 62.49 43.44 38.19 54.43 55.74 91.53 89.24 77.01 60.24 83.08 83.11
DAIL-SQL (SC) 63.03 45.81 43.06 55.93 57.20 91.53 90.13 75.29 62.65 83.56 -
TA-SQL 63.14 48.82 36.81 56.32 - 93.50 90.80 77.60 64.50 85.00 -
SuperSQL 66.92 46.67 43.75 58.60 60.62 94.35 91.26 83.33 68.67 87.04 85.92
Dubo-SQL - - - 59.71 66.01 - - - - - -
CodeS-15B CodeS 65.19 48.60 38.19 57.63 63.22 94.76 91.26 72.99 64.46 84.72 83.52
CodeS + KaSLA 68.32 50.75 38.19 60.17 64.52 96.37 89.46 76.44 63.86 84.82 84.46
StarCoder2 -15B TA-SL 59.89 45.38 34.72 53.13 60.51 95.56 92.38 78.16 63.25 85.98 85.03
TA-SL + KaSLA 66.59 49.46 38.89 58.80 63.56 94.76 92.15 77.59 66.87 86.27 85.47
DTS-SQL 63.03 46.02 34.72 55.22 64.17 91.94 90.58 78.74 66.87 85.11 85.49
DTS-SQL + KaSLA 63.89 50.11 39.58 57.43 65.20 94.76 90.81 80.46 68.07 86.36 85.81
CodeS 68.00 51.40 39.58 60.30 65.04 94.76 91.26 72.99 64.46 84.72 83.52
CodeS + KaSLA 67.68 54.84 42.36 61.41 66.87 94.76 91.03 78.74 66.27 85.88 84.46
Deepseek -V3 E-SQL 67.24 51.61 41.67 60.10 62.16 95.57 91.93 72.99 64.46 85.21 83.81
E-SQL + KaSLA 69.40 53.55 43.75 62.18 65.32 95.97 92.38 73.56 65.06 85.69 85.72
CHESS 67.89 52.90 46.53 61.34 65.28 95.57 92.16 73.56 65.06 85.50 85.33
CHESS + KaSLA 69.94 54.84 47.92 63.30 67.21 96.37 92.83 74.14 65.66 86.17 86.35
Table 1: The SQL generation performance of enhanced text-to-SQL models using KaSLA, which features 1.6 B parameters combining DeepSeek-coder-1.3B and RoBERTa-Large, is evaluated in terms of Execution Accuracy (EX) (%) and Valid Efficiency Score (VES) (%) on the BIRD-dev and Spider-dev datasets. We utilize KaSLA to substitute the original schema linking component in models that already incorporate schema linking, or to replace the full schema prompt for models that do not.
5.3 Experimental Results of Schema Linking (RQ2)
To answer the RQ2, we conduct the experiments and report the evaluation results of schema linking methods on BIRD-dev in Figure 2. We measured the performance of these schema linking methods in both table linking and column linking. The results indicate that, under the enhanced F1 score which specifically evaluates missing elements, only KaSLA achieves high accuracy in the challenging task of column linking. Schema linking models using LLMs often perform well in table linking but fall short in column linking due to the lack of additional optimizations for missing elements. This highlights the advantages of KaSLA’s framework and demonstrates how enhanced metrics can reveal the significant issue of element missing in current schema linking practices.
Schema linking method Table linking Column linking
DTS-SL (starcoder2-15b) 61.94 23.71
CodeS-SL (starcoder2-15b) 58.21 29.56
RSL-SQL (deepseek-v3) 67.11 17.20
KaSLA (deepseek-coder-1.3B) 81.92 62.78
Table 2: Schema linking performance with the proposed enhanced linking metric on BIRD-dev.
Ablation EX
CodeS + KaSLA 60.17
w/o Binary scoring model 56.06
w/o Probabilistic scoring model 57.82
w/o Hierarchical strategy 58.34
w/o upper limit 1 in Eq. 10 53.26
Table 3: Ablation studies of the key components of KaSLA. We conduct the experiments for CodeS + KaSLA on BIRD-dev and use CodeS with CodeS-15B as the text-to-SQL backbone.
Dataset Schema linking model Binary scoring function with deepseek-coder-1.3B Probabilistic scoring model with RoBERTa-Large Estimation and dynamic programming Text-to-SQL Total time EX
BIRD-dev Full Schema / / / 5.15 s 5.15 s 57.63
KaSLA 3.5 s 0.12 s $<$ 0.01 s 1.85 s 5.48 s 60.17
Table 4: Inference time cost per instance of each component in KaSLA using CodeS-15B as the text-to-SQL model.
Model SuperSQL CodeS CodeS + KaSLA CodeS + KaSLA (Transfer)
BIRD-dev 58.60 60.30 60.17 60.02
Spider-dev 87.40 84.72 84.82 84.56
Table 5: Execution Accuracy (EX) (%) of KaSLA trained on Spider-train but evaluated on BIRD-dev, and vice versa, for cross-scenario transfer ability evaluation.
5.4 Empirical studies (RQ3)
Ablation Study.
We presented the ablation study results for the key components in KaSLA in Table 3, the results concerning the choice of language model for the probabilistic scoring model are shown in Table 8 in Apendix and the scaling up performance of the scoring model are shown in Table 7 in Apendix, respectively.
As shown in Table 3: (i) Removing the binary scoring model results in a lack of confirmation for high-confidence elements, leading to the inclusion of more redundant ones. This underscores the binary model’s role in filtering out less relevant elements by ensuring that elements with high certainty are correctly identified. (ii) Omitting the probabilistic scoring model risks missing potentially relevant elements, resulting in gaps within schema linking. The probabilistic model is essential for assigning a non-zero probability to all elements, thus capturing those with potential relevance. (iii) The hierarchical strategy effectively reduces the number of candidate columns, thereby lowering complexity and improving efficiency. By structuring the linking process in stages, KaSLA can more precisely target relevant schema components. (iv) Without the upper limit set at 1, the model introduces disproportionate comparisons among similarly relevant items, potentially excluding some relevant elements. This constraint is crucial to maintain balanced relevance assessments, preventing the overshadowing of items that might otherwise contribute positively to SQL generation. These ablation study highlights the significance of key components within KaSLA.
Transferability.
We conducted two experiments to evaluate KaSLA’s performance in cross-scenario transfer. We trained the binary and probabilistic scoring models on the Spider training dataset and evaluated it on the BIRD dev dataset, and vice versa, to explore its cross-scenario transfer ability. Results are provided in Table 5. We can find that pre-training KaSLA on public datasets yields results that outperform the baselines and are only slightly lower than what domain-specific fine-tuning would achieve. The results show that KaSLA demonstrates strong cross-scenario transferability, highlighting its robustness and adaptability across different data domains.
Inference time cost.
We evaluated the inference time cost of integrating KaSLA into text-to-SQL frameworks in Table 4. By utilizing the lightweight LLM DeepSeek-coder-1.3B and significantly reducing the number of tokens in the input prompts through optimized schema linking, KaSLA achieves improved SQL generation accuracy with only a modest increase in inference time. This showcases KaSLA’s efficiency in balancing performance and computational demands.
6 Related Work
Schema linking is essential for aligning natural language queries with database components, playing a pivotal role in the overall text-to-SQL process. Recent approaches utilize large language models (LLMs) to enhance schema linkage through methods such as semantic retrieval, historical query integration, and bidirectional linking (Pourreza and Rafiei, 2024; Cao et al., 2024; Zhang et al., 2024). Techniques like DAIL-SQL and E-SQL refine linkage by leveraging LLMs’ contextual understanding and query enrichment capabilities (Gao et al., 2024; Caferoğlu and Ulusoy, 2024). Retrieval-based strategies focus on retrieve information from database to improve SQL generation (Pourreza and Rafiei, 2024). Chess (Talaei et al., 2024) and CodeS (Li et al., 2024b) utilize semantic retrieval to improve schema linking by accessing relevant database information. Despite these advancements, many state-of-the-art text-to-SQL models emphasize fine-tuning strategy. DTS-SQL (Pourreza and Rafiei, 2024) fine-tunes a local LLM to select relevant tables from user queries, while Solid-SQL (Liu et al., 2025) introduces a data augmentation method for enhancing schema linking. DELLM (Hong et al., 2024a) finetuned a specialized data expert LLM to provide additional knowledge for schema linking and SQL generation. Approaches like TA-SQL (Qu et al., 2024) use in-context learning to guide LLMs, generating dummy SQL for linked element abstraction. Similarly, RSL-SQL (Cao et al., 2024) employs a bidirectional approach for schema simplification. Meanwhile, E-SQL (Qu et al., 2024) and (Maamari et al., 2024) note that schema linking challenges persist, leading some models to use the entire schema directly as input for SQL generation.
Different with the previous research, our KaSLA utilize a knapsack optimization framework to avoid the relevant elements missing and the inclusion of redundant elements, optimizing schema linking to enhance SQL generation accuracy effectively.
7 Conclusion
This paper introduced the Knapsack Schema Linking Agent (KaSLA), a pioneering approach to overcoming schema linking challenges in text-to-SQL tasks. Based on knapsack optimization, KaSLA effectively prevent the missing of relevant elements and excess redundancy, thereby significantly enhancing SQL generation accuracy. Our introduction of an enhanced schema linking metric, sets a new standard for evaluating schema linking performance. KaSLA utilizes a hierarchical linking strategy, starting with optimal table linking and then proceeding to column linking within selected tables, effectively reducing the candidate search space. In each step, KaSLA employs a knapsack optimization strategy to link potentially relevant elements while considering a limited tolerance for potential redundancy. Extensive experiments on the Spider and BIRD benchmarks have demonstrated that KaSLA can significantly enhance the SQL generation performance of existing text-to-SQL models by substituting their current schema linking processes. These findings underscore KaSLA’s potential to revolutionize schema linking and advance the broader capabilities of text-to-SQL systems, offering a robust solution that paves the way for more precise and efficient database interactions.
8 Limitation
There are mainly two limitations of this work. First, although KaSLA demonstrates significant advancements in schema linking accuracy and efficiency, its performance traditionally depends on a comprehensive training dataset with detailed ground truth linking results. However, our experiments indicate that KaSLA possesses a degree of transferability, showing promise even in scenarios where training data is sparse or less diverse. This adaptability suggests that KaSLA can maintain reasonable effectiveness across different database environments, although performance might still be impacted in highly dynamic or rapidly evolving schema structures. Additionally, the reliance on a fine-tuning process of LLMs, integrating DeepSeek-coder-1.3B, could pose scalability challenges in resource-constrained settings. Despite these limitations, KaSLA remains a powerful tool for schema linking optimization in text-to-SQL tasks, and future work could focus on enhancing its adaptability and efficiency.
9 Ethics Statement
We confirm that we have fully complied with the ACL Ethics Policy in this study. All the datasets are publicly available and have been extensively used in research related to text-to-SQL.
Appendix A Implementation Details.
We implemented the proposed KaSLA and baselines under the following settings: For top-$K$ used in the redundancy tolerance estimation of KaSLA, we set it to 30. For method using fine-tuned LLM, We utilized LoRA (Hu et al., 2021) for parameter-efficient fine-tuning. The learning rate was initialized to $41e-$ 4 with a cosine decay, the batch size was set to 16, and the training epoch was 3.
Appendix B Input formats
We provided the input and output formats of the binary scoring function of KaSLA introduced in section 4.1 in Table 6.
Type Input format
Full schema table department, columns = [ department__\__id (int |||| primary key||||values: 1), name (text||||values: State, Treasury), ranking (int |||| values: 1, 2)] table head, columns = [ head__\__id (int |||| primary key |||| values: 1), name (text |||| values: Tiger Woods), age (real |||| values: 67.0)] table management, columns = [ department__\__id ( int |||| primary key |||| values: 2), head__\__id (int |||| primary key |||| values: 5), temporary__\__acting (text |||| values: Yes)] foreign keys: management.head__\__id = head.head__\__id, management.department__\__id = department.department__\__id
Question How many heads of the departments are older than 56? Generate SQL to solve the above question and list the relevant columns in each table
Type Output format
Simulated SQL SELECT count(*) FROM head WHERE age > 56
Schema linking results "department": null, "management": null, "head": [ "age" ]
Table 6: Input and output formats of the binary scoring model.
Appendix C The Dynamic Programming Approach used in KaSLA
We illustrate the details of the dynamic programming algorithm used in KaSLA in Algorithm 1.
Input: relevance score set ${\mathcal{R}}$, redundancy score set ${\mathcal{W}}$, the upper redundancy tolerance ${\mathcal{U}}$
Output: selection set $\hat{\mathcal{H}}$
$n\leftarrow|{\mathcal{R}}|$, $^{}^\hat{\mathcal{H}}\leftarrow\{\}$
[…]
$keep\leftarrow\text{array of }(n+1)\times({\mathcal{U}}+1)$
initialized to False
for $i\leftarrow 1$ to $n$ do
       for $w\leftarrow 0$ to ${\mathcal{U}}$ do
             if ${\mathcal{W}}[i-1]\leq w$ then
                   if $({\mathcal{R}}[i-1]+\mathcal{A}[i-1][w-{\mathcal{W}}[i-1]])>\mathcal{A}[i-1][w]$ then
                         $\mathcal{A}[i][w]\leftarrow{\mathcal{R}}[i-1]+\mathcal{A}[i-1][w-{\mathcal{W}}[i-1]]$
                         $keep[i][w]\leftarrow\text{True}$
                  else
                         $\mathcal{A}[i][w]\leftarrow\mathcal{A}[i-1][w]$
            else
                   $\mathcal{A}[i][w]\leftarrow\mathcal{A}[i-1][w]$
$k\leftarrow{\mathcal{U}}$  for $i\leftarrow n$ downto $1$ do
       if $keep[i][k]$ then
             ^.$\hat{\mathcal{H}}.\text{add}(i-1)$
             $k\leftarrow k-{\mathcal{W}}[i-1]$
return $\hat{\mathcal{H}}$
Algorithm 1 Dynamic Programming-based 0-1 Knapsack Problem Optimization
Appendix D Complexity Analysis
The efficiency of KaSLA stems from its hierarchical approach, dynamic programming and inherent parallelism. KaSLA’s time complexity is determined by two main phases: table linking and column linking within the linked tables. For a database with $n_{t}$ tables and a table capacity of $C_{t}$, the table linking phase has a time complexity of $O(n_{t}C_{t})$. KaSLA then links columns within the selected table. Let $n_{c}$ and $C_{c}$ represent the number of columns and the column capacity of a selected table, respectively. The time complexity of the column linking phase is $O(n_{c}C_{c})$. Consequently, the total time complexity of KaSLA is $O(n_{t}C_{t}+n_{t}(n_{c}C_{c}))$, indicating that the complexity of KaSLA scales relevant with the database dimensions.
Appendix E The Ablation Studies for Different Language Models
We present the ablation studies on the choice of language models for the probabilistic scoring model, as well as the performance scaling of various language models for the binary scoring model used in KaSLA. These studies are detailed in Table 8 and Table 7, respectively.
Dataset BERT -Base BERT -Large RoBERTa -Base RoBERTa -Large SGPT -1.3B
BIRD-dev 59.12 59.52 59.65 60.17 59.97
Table 7: Execution Accuracy (EX) (%) of the scaling up performance about the choice of language model for binary scoring model. We conduct the experiments for CodeS + KaSLA on BIRD-dev and use CodeS with CodeS-15B as the text-to-SQL backbone.
Dataset Deepseek-coder -1.3B Deepseek-coder -6.5B Deepseek-coder -7B
BIRD-dev 60.17 61.28 61.47
Table 8: Execution Accuracy (EX) (%) of the ablation study about the choice of language model for probabilistic scoring model. We conduct the experiments for CodeS + KaSLA on BIRD-dev and use CodeS with CodeS-15B as the text-to-SQL backbone.